# TeViR: Text-to-Video Reward with Diffusion Models for Efficient Reinforcement Learning

### v1

@@ Abstract
Developing scalable and generalizable reward engineering for reinforcement learning (RL) is crucial for creating general-purpose agents, especially in the challenging domain of robotic manipulation. While recent advances in reward engineering with Vision-Language Models (VLMs) have shown promise, their sparse reward nature significantly limits sample efficiency. This paper introduces TeViR, a novel method that leverages a pre-trained text-to-video diffusion model to generate dense rewards by comparing the predicted image sequence with current observations. Experimental results across 11 complex robotic tasks demonstrate that TeViR outperforms traditional methods leveraging sparse rewards and other state-of-the-art (SOTA) methods, achieving better sample efficiency and performance without ground truth environmental rewards. TeViR’s ability to efficiently guide agents in complex environments highlights its potential to advance reinforcement learning applications in robotic manipulation.
@@ I Introduction
Although inverse reinforcement learning (IRL) [9] learns rewards from pre-collected expert demonstration, these learned reward functions are unreliable for learning policies due to noise and misspecification errors [10], especially for robotic manipulation tasks since in-domain data is limited [11, 12]. Additionally, the learned reward functions is not generally applicable across tasks. With the impressive generalization capabilities, pre-trained Vision-Language Models (VLMs) enables referencing visual features with natural language [13], and thus used to generate sparse rewards for policy training by calculating the similarity between language descriptions of tasks and the visual observations [14, 15, 11]. However, the reward signals generated in these methods are often unstable and meets high variance, leading to insufficient supervision and low sample efficiency. This is because VLMs lack the diversity to generalize to different tasks as they are trained only on matching language descriptions and visual observations [16].
@@
To this end, we present Text-to-Video Reward (TeViR), a novel framework designed to calculate dense rewards by generating videos with the language instruction input, enabling agents to efficiently learn the desired policy. Our framework involves video generation for planning the task execution and assessing whether current observation meet expectations and contribute to task progress. Visual and linguistic information play crucial roles in human learning and decision-making processes [27, 19, 28]. We combine these two types of information with prior information across different tasks, and utilizes text-to-video diffusion models for long horizon planning, which generates observations of the imaginary trajectory based on the current visual observation and linguistic task description. By comparing these predicted images with the current observations, dense rewards can be calculated for each timestep, without relying on the environment rewards. Such dense rewards enable the policy to evaluate the current state and efficiently learn the desired behavior [29, 30]. Benefits from the long-term planning ability of the video generation model, TeViR exploits the information from generated trajectory, thereby achieving better performance and sample efficiency compared to predicting the short-term observations.
@@
    We conduct experiments to show that TeViR achieves the new SOTA for the reward engineering in 11 Meta-World tasks. Notably, without the environmental rewards, the average success rate on 8 robot manipulation tasks increased by 49.1% within 2M training steps.
@@
Figure 1: Overall framework of TeViR. The text-to-video diffusion model takes the initial RGB observation of the environment and a textual goal description as its input. Then it synthesizes an image sequence of the future and the TeViR is calculated using the current observation and the synthesized image sequence. Finally, the TeViR will help the agent update its policy.
@@ II-A Learning from Expert Videos
How to learn policy from expert videos has been studied for decades. Imitation learning (IL) methods often rely on state-action pairs from expert demonstrations to learn policy. Recent studies such as [31, 32, 33], emphasize learning directly from the raw export data, bypassing the need for manually labeled data. IRL [34, 35, 36] extends the capabilities of IL by deriving reward functions from the observed behaviors of experts. Several works use perceptual metrics such as goal-image similarity and trajectory similarity to calculate rewards [37, 38, 39, 12, 40]. A novel approach involves using video prediction models to estimate future states and derive rewards based on predicted log-likelihood or entropy [14, 24, 25]. These methods allow for dynamic reward generation, potentially increasing the adaptability and efficiency of learning algorithms. In contrast to these methods, our approach leverages a text-to-video diffusion model to generate dense rewards by comparing predicted video sequences with current observations, offering a more scalable and generalizable reward signal for RL.
@@ II-B Large models as Reward Functions
To design reward functions capturing human preferences, methods [8, 41] explores large language models (LLMs) in text-based games like negotiations or card games. Other works followed by demonstrating that LLMs can write code of the reward function for training robots [42, 43, 44]. Focusing on applications that interact with the physical world, approaches employ VLMs as a success detector to provide sparse task rewards [45], or to evaluate whether image observation of agent aligns with language description of the robot tasks [46, 15, 47, 11, 48]. However, the accuracy of these generated rewards relies heavily on the input task specification and how well-aligned the agent visual observations are to the dataset used for VLMs pre-training [15, 12].
Other than using language as the interface, more recent approaches applied the large Text-to-Image model to generate visual goal [40] or web-scale text-to-video diffusion model to learn universal policies [19, 20, 49], serving image as a unified representation of the physical world. In contrast, our method learns dense rewards via a text-to-video diffusion model conditioned on both language descriptions and current visual observations, offering more informative and continuous feedback to accelerate policy learning and enhancing the scalability and generalization across various tasks.
@@ IV-A TeViR Framework
As depicted in Figure 1, our framework leverages expert experience from pre-trained text-to-video diffusion models at a high level. The model takes a language description to specify desired task and an RGB image to indicate the initial state of the environment as condition, and generates the future image sequence til task completion. We expect the RGB image to capture dynamic features in the environment, with the generated expert video bridging the domain gap between task language description and the robot behavior. This enables more comprehensive guidance for long-horizon tasks, addressing the challenges posed by sparse or unreliable reward signals.
@@ IV-B Multi-view Video Generation
Figure 3: Progress calculation illustration. For each image $z_{t}$ in a trajectory, the progress reward will calculate the similarity with the corresponding generated image sequence within its progress $M_{t}$. We assume that the image $z_{t-1}$ on a trajectory with progress $M_{t-1}=4$ is similar to the farthest reached image $z_{M_{t-1}}$ and we have $M_{t}=5$. If $z_{t}$ of the trajectory is similar to image $z_{M_{t}}$, then progress for $z_{t+1}$ should be $6$, labeled in green. Otherwise, the progress for $z_{t+1}$ would maintain $5$, labeled in orange.
@@ IV-C Text-to-video Reward
We then have the progress reward $r^{prog}_{t}$, for evaluating the overall progress toward completing the task. Consider the task of opening a door. If the agent skips the step of grasping the doorknob and tries to pull the door directly, it will fail to complete the task correctly. Therefore, we should guide the agent to perform each critical step in sequence to ensure the proper execution of the task.
@@ V Experiments
    What implementation detail of TeViR matters for task learning? (Section V-E)
@@ V-A1 Baselines
  • RoboCLIP: RoboCLIP [^3] https://github.com/sumedh7/RoboCLIP [15] utilizes pretrained VLMs to generate rewards for the agent. This is done by providing a sparse reward to the agent at the end of the trajectory which describes the similarity of the agent’s behavior to the linguistic task description. This comparison tests the benefit of generating dense rewards for policy learning.
@@
TABLE I: Linguistic description and hyperparameters of TeViR for each task
Task Linguistic Description $\theta$$w^{left},w^{top},w^{close}$}
@@
Simulation Environments and Training Details
We intend to demonstrate the effectiveness of TeViR on visual robotic manipulation tasks from Meta-World [53], with a Sawyer robot arm. Each task is associated with $64\times 64$-dimensional RGB images of three view and a binary ground truth environmental sparse reward. To ensure task diversity and complexity when evaluating reward function, we choose 11 manipulation tasks, including different skills such as reaching, pushing, and grasping. These tasks are widely used in visual RL and are chosen to be diverse in objects and manipulating skills. The linguistic description and hyperparameters of TeViR of each task is shown in Table I.
We collect 30 expert videos of 3 view for each via the scripted policy provided by the official repository [^7] https://github.com/Farama-Foundation/Metaworld. And we utilize the task description provided by the Meta-World benchmark to synthesize ten sentences with consistent semantic content to expand data diversity.
For the video generation, we used a fixed pre-trained CLIP-Text encoder (63M parameters) [54] to encode linguistic task descriptions into 512-dimension embedding. Then we follow AVDC [20] and use a similar network architecture and training setup which utilizes Video U-Net model with 3 residual blocks of 128 base channels and channel multiplier [1, 2, 3, 4, 5], attention resolutions [8, 16], attention head dimension 32. We use 1e-4 as the learning rate and Adam optimizer for model updating. The training is performed on an NVIDIA A6000 GPU with a batch size of 8, using the first 80% of the video data, with the remaining 20% used for evaluation. The video diffusion model is trained for 400K steps, taking approximately 20 hours to complete.
Note that for a fair comparison, all reward engineering methods tested are agnostic to the choice of downstream RL algorithm. We use DrQv2888https://github.com/facebookresearch/drqv2 [55] as the backbone. The discount is 0.99. And we use $action\_repeat=3$ and $feature\_dim=50$ for DrQv2 policy. We use 1e-4 as the learning rate and Adam optimizer for policy training, which is performed on an NVIDIA A6000 GPU with a batch size of 256.
Figure 5: Performances of the different reward engineering methods with the encironmental sparse reward on 8 tasks from Meta-World benchmark. Raw Sparse Reward and Diffusion Reward are compared on each task with 3 random seeds within 1M training steps. Our method achieves better performance by incorporating the ground truth sparse reward.
@@
Figure 6: Performances of the different reward engineering methods without the environmental feedback on 8 tasks from Meta-World benchmark. RoboCLIP, VIPER and Diffusion Reward- are compared on each task with 3 random seeds within 2M training steps. Our method achieves best performance on all tasks without the environmental sparse reward.
@@ V-C Learning without Ground Truth Sparse Reward
The experimental results demonstrate that our method achieves outstanding performance across all tasks in the absence of the ground truth environmental sparse rewards. This finding underscores the robustness and effectiveness of our approach in complex robot manipulation tasks. As depicted in Figure 4, RoboCLIP, which generates a sparse reward by calculating the similarity between the image trajectory and the linguistic description of the desired behavior, fails in almost all tasks. The agent can only get closer to its goal position, as language alone is insufficient to capture the intricacies of complex behaviors.
V-D Learning with Poorly Generated Video
To simulate poor-quality video generation, we will employ two distinct methods. These evaluations aim to provide a comprehensive understanding of the robustness of our method and UniPi under adverse video generation conditions. All results for UniPi are implemented via open-loop control and tested on 30 independent experiment with random seeds. All results for TeViR are reported by 3 random seeds within 2M training steps.
@@ V-D Learning with Poorly Generated Video
Impact of Added Noise
The first method involves directly adding noise to the generated videos. We will investigate the performance of the two methods under varying signal-to-noise ratios (SNR). The results are shown in Table II, and we provide training curves of TeViR in Figure 8.
Impact of Task-Irrelevant and Temporally Disordered Frames
The second method introduces task-irrelevant or temporally disordered frames into the generated videos. We will explore the final performance of both methods under a fixed proportion of erroneous frames. The results are shown in Table III, and we provide training curves of TeViR in Figure 8.
@@
V-E Ablation Studies
V-E1 Multi-view Observations
Selecting observations from multiple view is crucial for enhancing the environment representation, which directly impacts reward calculation, as evidenced in Figure 9. The $close$ view primarily provides detailed information about the manipulated object, making it insufficient for guiding the agent effectively on its own. In contrast, training without the $close$ view may lose essential information about the manipulated object, which leads to unstable training, especially for those tasks that require careful manipulation on the object (e.g. ’window-close-v2’). On the other hand, the $left$ and $top$ view offer comprehensive information about the environment and the robot arm, leading to better performance when they are used. The weights assigned to each view can be adjusted to meet the specific requirements of different tasks and environments, as shown in Table I. For instance, the ’window-close-v2’ task necessitates a parallel movement to close the window, thus requiring a higher weight on the $top$ view. Conversely, the ’drawer-open-v2’ task demands precise manipulation of the handle, necessitating a higher weight on the $close$ view. Implementing multiple view can improve the accuracy of similarity calculations, offering more efficient and stable guidance, and thereby achieving better overall performance.
@@ V-E1 Multi-view Observations
V-E2 Component Analysis of TeViR

### v2

@@ Abstract
Developing scalable and generalizable reward engineering for reinforcement learning (RL) is crucial for creating general-purpose agents, especially in the challenging domain of robotic manipulation. While recent advances in reward engineering with Vision-Language Models (VLMs) have shown promise, their sparse reward nature significantly limits sample efficiency. This paper introduces TeViR, a novel method that leverages a pre-trained text-to-video diffusion model to generate dense rewards by comparing the predicted image sequence with current observations. Experimental results across 13 simulation and real-world robotic tasks demonstrate that TeViR outperforms traditional methods leveraging sparse rewards and other state-of-the-art (SOTA) methods, achieving better sample efficiency and performance without ground truth environmental rewards. TeViR’s ability to efficiently guide agents in complex environments highlights its potential to advance reinforcement learning applications in robotic manipulation.
@@ I Introduction
Although inverse reinforcement learning (IRL) [9] learns rewards from pre-collected expert demonstration, these learned reward functions are unreliable for learning policies due to noise and misspecification errors [10], especially for robotic manipulation tasks since in-domain data is limited [11]. Additionally, the learned reward functions is not generally applicable across tasks. With the impressive generalization capabilities, pre-trained Vision-Language Models (VLMs) enables referencing visual features with natural language [12], and thus used to generate sparse rewards for policy training by calculating the similarity between language descriptions of tasks and the visual observations [13, 14]. However, the reward signals generated in these methods are often unstable and meets high variance, leading to insufficient supervision and low sample efficiency. This is because VLMs lack the diversity to generalize to different tasks as they are trained only on matching language descriptions and visual observations [15].
@@
To this end, we present Text-to-Video Reward (TeViR), a novel framework designed to calculate dense rewards by generating videos with the language instruction input, enabling agents to efficiently learn the desired policy. Our framework involves video generation for planning the task execution and assessing whether current observation meet expectations and contribute to task progress. Visual and linguistic information play crucial roles in human learning and decision-making processes [26, 18, 27]. We combine these two types of information with prior information across different tasks, and utilizes text-to-video diffusion models for long horizon planning, which generates observations of the imaginary trajectory based on the current visual observation and language task description. By comparing these predicted images with the current observations, dense rewards can be calculated for each timestep, without relying on the environment rewards. Such dense rewards enable the policy to evaluate the current state and efficiently learn the desired behavior [28, 29]. Benefits from the long-term planning ability of the video generation model, TeViR exploits the information from generated trajectory, thereby achieving better performance and sample efficiency compared to predicting the short-term observations.
@@
    We conduct experiments to show that TeViR achieves the new SOTA for the reward engineering in 11 Meta-World tasks and 2 real-world tasks. Notably, without the environmental rewards, the average success rate increased by 49.1% on Meta-World tasks and by 29.0% on real-world tasks within same training steps.
@@
Figure 1: Overall framework of TeViR. The text-to-video diffusion model takes the initial RGB observation of the environment and a textual goal description as its input. Then it synthesizes an image sequence of the future and the TeViR is calculated using the current observation and the synthesized image sequence. Finally, the TeViR helps the agent update its policy.
@@ II-A Learning from Expert Videos
How to learn policy from expert videos has been studied for decades. Imitation learning (IL) methods often rely on state-action pairs from expert demonstrations to learn policy. Recent studies such as [30, 31, 32], emphasize learning directly from the raw export data, bypassing the need for manually labeled data. IRL [33, 34, 35] extends the capabilities of IL by deriving reward functions from the observed behaviors of experts. Several works use perceptual metrics such as goal-image similarity and trajectory similarity to calculate rewards [36, 37, 11, 38]. A novel approach involves using video prediction models to estimate future states and derive rewards based on predicted log-likelihood or entropy [13, 23, 24]. These methods allow for dynamic reward generation, potentially increasing the adaptability and efficiency of learning algorithms. In contrast to prior methods, our approach utilizes a text-to-video diffusion model to generate dense rewards by comparing predicted video sequences with the agent’s current observations, providing a more scalable and generalizable feedback mechanism for RL.
@@ II-B Large models as Reward Functions
To design reward functions capturing human preferences, methods [8, 39] explores large language models (LLMs) in text-based games like negotiations or card games. Other works followed by demonstrating that LLMs can write code of the reward function for training robots [40, 41, 42]. Focusing on applications that interact with the physical world, approaches employ VLMs as a success detector to provide sparse task rewards [43], or to evaluate whether image observation of agent aligns with language description of the robot tasks [44, 14, 45, 46, 47]. However, these methods operate on individual frames or short clips, without modeling full temporal task completion, and their reward accuracy heavily depends on how well the agent’s visual observations align with the data distribution of the pre-trained VLMs[14, 11]. This limits their effectiveness in long-horizon, open-ended tasks, particularly when domain or viewpoint shifts occur.
Other than using language as the interface, more recent approaches applied the large Text-to-Image model to generate visual goal [38] or web-scale text-to-video diffusion model to learn universal policies [18, 19, 48], serving image as a unified representation of the physical world. In contrast, our method learns dense rewards via a text-to-video diffusion model conditioned on both language descriptions and current visual observations, offering more informative and continuous feedback to accelerate policy learning while improving the scalability and generalization across various tasks.
@@ IV-A TeViR Framework
As depicted in Figure 1, our framework leverages expert experience from pre-trained text-to-video diffusion models at a high level. The model takes a language description to specify desired task and an RGB image to indicate the initial state of the environment as condition, and generates a sequence of key frames illustrating the expected progression toward task completion. We expect the RGB image to capture dynamic features in the environment, with the generated expert video bridging the domain gap between task language description and the robot behavior. This enables more comprehensive guidance for long-horizon tasks, addressing the challenges posed by sparse or unreliable reward signals.
@@ IV-B Multi-view Video Generation
Figure 3: Progress calculation illustration. For each image $z_{t}$ in a trajectory, the progress reward calculates the similarity with the corresponding generated image sequence within its progress $M_{t}$. We assume that the image $z_{t-1}$ on a trajectory with progress $M_{t-1}=4$ is similar to the farthest reached image $z_{M_{t-1}}$ and we have $M_{t}=5$. If $z_{t}$ of the trajectory is similar to image $z_{M_{t}}$, then progress for $z_{t+1}$ should be $6$, labeled in green. Otherwise, the progress for $z_{t+1}$ would maintain $5$, labeled in orange.
@@ IV-C Text-to-video Reward
We then have the progress reward $r^{prog}_{t}$, for evaluating the overall progress toward completing the task. Consider the task of opening a door. If the agent skips the step of grasping the doorknob and tries to pull the door directly, it fails to complete the task correctly. Therefore, we should guide the agent to perform each critical step in sequence to ensure the proper execution of the task.
@@ V Experiments
    What implementation detail of TeViR matters for task learning? (Section V-F)
@@ V-A1 Baselines
  • RoboCLIP: RoboCLIP [^3] https://github.com/sumedh7/RoboCLIP [14] utilizes pre-trained VLMs to generate rewards for the agent. This is done by providing a sparse reward to the agent at the end of the trajectory which describes the similarity of the agent’s behavior to the language task description. This comparison tests the benefit of generating dense rewards for policy learning.
@@
TABLE I: Language description and hyperparameters of TeViR for each task
Task Language Description $\theta$$w^{left},w^{top},w^{close}$}
@@
Pick Carrot (real-world) Put the carrot into the plate. 0.7 {0.8, 0, 0.5}
Pull Compartment (real-world) Pull open the sliding compartment. 0.7 {0.8, 0, 0.5}
@@
Figure 5: Performances of the different reward engineering methods with the environmental sparse reward on 8 tasks from Meta-World benchmark. Raw Sparse Reward and Diffusion Reward are compared with TeViR+ with 3 random seeds within 1M training steps.
Experimental Environments and Training Details
We intend to demonstrate the effectiveness of TeViR on visual robotic manipulation tasks from Meta-World [52] with Sawyer arm, and real-world tasks with Franka Emika arm. To ensure task diversity and complexity when evaluating reward function, we choose 11 Meta-World manipulation tasks and 2 real-world tasks shown in Figure 6, including different skills such as reaching, pushing, and grasping. Each task is associated with $64\times 64$-dimensional RGB images of multiple view and a binary ground truth environmental sparse reward. The sparse reward in real-world tasks is generated by a pre-trained ResNet-based classifier, which evaluates the completion of target task. These tasks are widely used in visual RL and are chosen to be diverse in objects and manipulating skills. The language description and hyperparameters for each task are shown in Table I.
Figure 6: Overview of all real-world experimental tasks, including a) Pick Carrot and b) Open Drawer.
For the Meta-World tasks, we collect 30 expert videos of each view for each via the scripted policy provided by the official repository [^7] https://github.com/Farama-Foundation/Metaworld. For the real-world tasks, we collect 30 expert videos of each view through human tele-operation. And we utilize the task description as shown in Table I to synthesize ten sentences with consistent semantic content to expand data diversity.
For the video generation, we used a fixed pre-trained CLIP-Text encoder (63M parameters) [53] to encode language task descriptions into 512-dimension embedding. Then we follow AVDC [19] and use a similar network architecture and training setup which utilizes Video U-Net model with 3 residual blocks of 128 base channels and channel multiplier [1, 2, 3, 4, 5], attention resolutions [8, 16], attention head dimension 32. We use 1e-4 as the learning rate and Adam optimizer for model updating. The training is performed on an NVIDIA A6000 GPU with a batch size of 8, using the first 80% of the video data, with the remaining 20% used for evaluation. The video diffusion model is trained for 400K steps, taking approximately 20 hours to complete.
Note that for a fair comparison, all reward engineering methods tested are agnostic to the choice of downstream RL algorithm. For Meta-World tasks, we use DrQv2888https://github.com/facebookresearch/drqv2 [54] as the backbone. And we use $action\_repeat=3$ and $feature\_dim=50$ for DrQv2 policy. For real-world tasks, we adopt HIL-SERL [^9] https://github.com/rail-berkeley/hil-serl [55], incorporating human intervention during training. For all tasks, the discount is 0.99. We use 1e-4 as the learning rate and Adam optimizer for policy training, which is performed on an NVIDIA A6000 GPU with a batch size of 256.
Figure 7: Performances of the different reward engineering methods without the environmental feedback on 8 tasks from Meta-World benchmark. RoboCLIP, VIPER and Diffusion Reward- are compared with TeViR with 3 random seeds within 2M training steps.
@@ V-C Learning without Ground Truth Sparse Reward
The experimental results demonstrate that our method achieves outstanding performance across all tasks in the absence of the ground truth environmental sparse rewards. This finding underscores the robustness and effectiveness of our approach in complex robot manipulation tasks. As depicted in Figure 4, RoboCLIP, which generates a sparse reward by calculating the similarity between the image trajectory and the language description of the desired behavior, fails in almost all tasks. The agent can only get closer to its goal position, as language alone is insufficient to capture the intricacies of complex behaviors.
@@
V-D Learning with Poorly Generated Video
To simulate poor-quality video generation, we employ two distinct methods. These evaluations aim to provide a comprehensive understanding of the robustness of our method and UniPi under adverse video generation conditions. All results for UniPi are implemented via open-loop control and tested on 30 independent experiment with random seeds. All results for TeViR are reported by 3 random seeds within 2M training steps.
Impact of Added Noise
The first method involves directly adding noise to the generated videos. We investigate the performance of the two methods under varying signal-to-noise ratios (SNR). The results are shown in Table II, and we provide training curves of TeViR in Figure 9.
Impact of Task-Irrelevant and Temporally Disordered Frames
The second method introduces task-irrelevant or temporally disordered frames into the generated videos. We explore the final performance of both methods under a fixed proportion of erroneous frames. The results are shown in Table III, and we provide training curves of TeViR in Figure 9.
@@ V-D Learning with Poorly Generated Video
Figure 10: Trajectory returns of TeViR on real-world ”Open Drawer” task. Only Traj. 3 successfully completes the task.
Figure 11: Performances Raw sparse reward and TeViR+ on 2 real-world tasks. Raw Sparse Reward is compared with TeViR+ and all results are reported as a running average over 20 episode.
V-E Real World Experiments
To further assess the practical applicability of TeViR in real-world settings, where reward design remains a fundamental challenge, we extend our experiments to a physical robotic platform. Specifically, we evaluate RL training using TeViR+ against a baseline that relies solely on raw sparse rewards across two representative manipulation tasks, under otherwise identical training conditions.
As shown in Figure 10, we first validate the effectiveness of TeViR rewards in a real-world ”Open Drawer” task. In this setup, three trajectories reflect varying levels of task progress: Traj.1 exhibits only minor movement, Traj.2 fails to grasp the drawer handle, and only Traj.3 completes the task successfully. TeViR accurately distinguishes these differences, assigning rewards that reflect task success without relying on any hand-crafted metrics.
To fairly isolate the impact of TeViR+, we ensure a consistent level of human intervention across all settings. As shown in Figure 11, TeViR+ significantly accelerates policy learning and leads to more stable and goal-consistent behaviors in real-world conditions. Notably, it achieves higher success rates within the same training steps, outperforming the sparse-reward baseline (e.g., 100% vs. 65% and 90% in the two tasks). These results highlight TeViR’s potential to reduce the reliance on hand-crafted reward functions while preserving sample efficiency and robustness.
@@ V-E Real World Experiments
V-F Ablation Studies
V-F1 Multi-view Observations
Selecting observations from multiple view is crucial for enhancing the environment representation, which directly impacts reward calculation, as evidenced in Figure 12. In our Meta-World experiments, camera viewpoints are fixed, and we adjust a small number of hyperparameters, such as the weights assigned to each view, to align with the specific requirements of each task. The $close$ view primarily provides detailed information about the manipulated object, making it insufficient for guiding the agent effectively on its own. In contrast, training without the $close$ view may lose essential information about the manipulated object, which leads to unstable training, especially for those tasks that require careful manipulation on the object (e.g. ’window-close-v2’). On the other hand, the $left$ and $top$ view offer comprehensive information about the environment and the robot arm, leading to better performance when they are used. The weights assigned to each view can be adjusted to meet the specific requirements of different tasks and environments, as shown in Table I. For instance, the ’window-close-v2’ task necessitates a parallel movement to close the window, thus requiring a higher weight on the $top$ view. Conversely, the ’drawer-open-v2’ task demands precise manipulation of the handle, necessitating a higher weight on the $close$ view. Implementing multiple view can improve the accuracy of similarity calculations, offering more efficient and stable guidance, and thereby achieving better overall performance. While in real-world settings, we can place cameras to adequately capture the environment, the manipulated objects, and the robot arm’s state, while minimizing occlusions, without the need for additional hyperparameter tuning as shown in Figure 10. This physical flexibility preserves generality across tasks without requiring manual adjustment of view weights, while still ensuring effective observations for reward computation and policy learning.
V-F2 Component Analysis of TeViR